# Trojan Cleansing with Neural Collapse

### v1

@@ 1 Introduction
Over the past decade, deep neural networks have achieved state-of-the-art performance in a vast array of sensitive machine learning (ML) tasks such as autonomous driving, medical diagnostics, and financial portfolio management. And with the advent transformer architectures and large language models, these ML technologies are being incorporated in an ever increasing number of applications and products. The unprecedented speed of this adoption, along with prior research establishing the vulnerability of these methods to a wide variety of attacks, has made the study of security and safety of our ML models an important area of research. This paper considers an important class of training-time attacks on these models known as trojan attacks. In particular, our work attempts to 1) draw upon new results in deep learning theory to understand how trojan attacks alter the structure of neural networks, 2) experimentally validate these conclusions, and 3) use these insights to design and test a lightweight, broadly applicable method for mitigating trojan attacks.
Trojan attacks, or backdoor attacks, aim to maliciously taint the neural network training process so that the trained model so that the presence of the trojan trigger on any input will force the network to produce a target classification [10]. This type of attack is generally accomplished by means of data poisoning, where the adversary adds their chosen trigger to a small portion of a training dataset and labels the poisoned samples with their selected target class. Models trained on this dataset then learn both the correct mappings for clean data as well as the malicious association for inputs containing the trigger. Figure 1 illustrates an example of clean data point (a) along a poisoned one (b). While the yellow trigger in this example is quite visible, as shown by [24], poisoning may take on a subtle nature that is undetectable by human auditors so that that malicious behavior can be hidden until test time or model deployment. This, alongside the incredible amounts of data scraped for use in training very large models, makes trojans a tangible threat even for users training their own models.
Figure 1: An example of trojan attack on a traffic sign classifier. (a) model works normally on clean data without triggers, (b) attackers manipulate the prediction by adding a specific trigger. Image credit: [26].
In this paper, for the first time, we make a concrete connection between training with poisoned data and a phenomenon recently studied in the deep learning theory literature, called Neural Collapse (NC) [22]. Neural Collapse describes the widely observed phenomenon wherein, over the course of training a neural network, highly symmetric structures emerge in both the network’s weights and the feature representations of training samples. In contrast, trojan attacks are inherently asymmetric: they require any input be classified to a single target class given only the presence of a small perturbation. This tension motivates our work here.
@@ 1.1 Contributions
While there is a large body of work on trojan detection and mitigation, there has been limited investigation into the mechanism by which trojan triggers act within neural networks with the exception of the topological analysis in [33] observing that the trojaned networks tend to have “skip connections” identified by abnormally strong correlation between neurons in early and late layers in the network. Our analysis is similar in nature as we try to explain and address the problem through an analytic lense.
We pursue this question instead through the inherent asymmetry introduced by training on poisoned datasets and the natural impact this has on the Neural Collapse phenomenon. We discuss NC in depth in Section 2.3, but, briefly, NC is a behavior of over-parameterized neural networks that are trained for an extended period of time after reaching perfect training accuracy. This behavior, associated with generalization [23] and robustness [22], ensures that, when trained on clean datasets, the final feature representations of the training data and the weight matrix of the final layer of the network both converge to a simple, highly symmetric geometric structure. This symmetric structure, however, comes directly in contradiction to the asymmetry introduced by data poisoning. In this work, we present extensive experimental evidence that the process of embedding a trojan in a neural network significantly weakens the NC phenomenon in measurable ways.
We leverage these insights to initiate a new class of trojan cleansing methods: algorithms which take a potentially trojaned network and remove the trojan trigger while maintaining the network’s performance on clean data. Many existing cleansing methods, as detailed in Section 2.2, rely either on some form of model compression, which may harm the model’s generalization capabilities, or on reconstructing the trojan trigger, which relies on knowledge of the class of the triggers the adversary may have chosen from. We are able to instead make a small, carefully targeted adjustment to the weights of the network without any knowledge of the possible trojan triggers.
@@ 2 Related Work
Figure 2: Visualization of NC. The figure illustrates NC as training progresses from left to right. Green spheres represent the vertices of the standard ETF, red spheres and connecting lines represent the linear classifiers, blue spheres and lines represent class means, and small blue spheres represent last-layer features. Each class is distinguished by different color shade. As training progresses, the last-layer features converge onto their respective class means, the class means align with the vertices of the simplex ETF, and the linear classifiers approach (the dual of) their corresponding class means, figure due to [22]
@@ 2.2 Model Cleansing
Many approaches to model cleansing task use model compression techniques to remove the portion of the network responsible for the action of the trigger. Li et al. [15] and Xia et al. [29] use forms of distillation to preserve only the model’s performance on clean data while Wang et al. [27] identifies and prunes neurons likely associated with the trojan trigger. Other approaches rely on first reconstructing potential triggers: [4] learns a generative model to create likely triggers and then corrects the network via adversarial training. Zeng et al. [31] proposes a minimax formulation and computationally efficient optimization scheme for trigger reconstruction. Other approaches include [13], which leverages sparse reconstruction of inputs and feature representations to remove potential triggers, and [35], which applies adversarial training to a new layer inserted into the network. [34] is a detection and mitigation method that shrinks the norms of backdoor-related neurons by using sharpness-aware minimization with fine-tuning. Wu et al. [28] introduced BackdoorBench, a platform which implements many of these algorithms.
Related, but distinct work focuses only on detecting whether or not a model may be trojaned. A small selection of such works include those focusd on data analysis, as in Baracaldo et al. [2], Liu et al. [16], Shen et al. [25] and those which directly analyze the model, such as Chen et al. [3], Ma et al. [19], Fields et al. [9].
@@ 2.3 Neural Collapse
Neural Collapse is a phenomenon, first reported by Papyan et al. [22] and thoroughly reviewed by [14], which has been observed across a wide variety datasets and model architectures. It describes a process observed in over-parameterized classification deep neural networks which are trained well past achieving zero training error, into a period referred to as the terminal phase of training (TPT). As depicted in Figure 2, as the TPT progresses, the final feature representations of the training data and the weight matrix of the network’s last layer gradually converge to a highly uniform, symmetric geometric structure known as a simplex equiangular tight frame (simplex ETF).
While NC is associated with models trained well past perfect training accuracy, this practice is known to have practical benefits as in the double descent phenomenon [20]. Further experiments in [22] demonstrate in particular that the TPT offers several practical advantages including improved generalization performance, enhanced robustness, and increased interpretability. Various theoretical models, based on locally elastic dynamics [32] and unconstrained features [36, 35] have begun examining how NC arises and its connections to these other neural network characteristics.
It is important to note that Neural Collapse is not a binary phenomenon that either does or does not occur. In practice the characteristics of collapse gradually arise to varying extents over the course of the TPT. Consequently, our discussion of NC will focus on its degree, quantified by four metrics which converge to 0 as the various qualitative characteristics of NC are attained [22]. We will describe these metrics in detail in Section 3.2 after introducing notation.
3 Methodology
3.1 Problem Setting
@@
The trojan attacks we evaluate in our experiments are carried out via data poisoning. To implement such an attack, we first choose a trigger function $\kappa:\mathbb{R}^{d}\to\mathbb{R}^{d}$. In our experiments this function will be the overlay of a small, solid patch on the base image, as shown in Figure 1. We then choose a poisoning proportion $\delta\in[0,1)$ and a target class $k_{T}\in[K]$. For each $\bm{X}_{k}$ with $k\neq k_{T}$ and each sample $\bm{x}\in\bm{X}_{k}$, with probability $\delta$ we delete $\bm{x}$ from $\bm{X}_{k}$ and add $\kappa(\bm{x})$ to $\bm{X}_{k_{T}}$. We say a model is trojaned if $\delta>0$, and a model is benign if $\delta=0$.
@@
3.2 Neural Collapse Metrics
The observed phenomenon of Neural Collapse can be more precisely described by a set of four characteristics laid out by [22] and quantified by metrics defined in [11]. We will use these quantities to track the degree of NC occurring over training for both benign and trojaned models, and we will show that embedding a trojan trigger in a model significantly slows and weakens NC along multiple dimensions.
NC1: Variability collapse The final feature representation of every training sample in every class converges to the mean of the feature representations of the samples from their respective classes. Equivalently, the variability of feature representations across any given class goes to 0. This can be quantified by defining the within-class and between-class covariances
$\Sigma_{B}$ $\coloneqq\frac{1}{K}\sum_{k=1}^{K}\bm{m}_{k}\bm{m}_{k}^{T}.$
We can then quantify the degree of variability collapse by the relationship between the spread of features within a class relative to the spread between the feature means of different classes:
[…]
NC2: Convergence to simplex ETF The rows of both $\bm{M}$ and $\bm{W}$ converge to a simplex ETF. In particular, this means that the norms of each of the rows equalize and that they become maximally pairwise separated from each other as measured by cosine similarity. Note that the maximum pairwise cosine similarity among a set of $K$ vectors is $-(K-1)^{-1}$. These two characteristics can be quantified, respectively, by the two statistics
[…]
We note that this defines the NC2 metric for both the centered class means, $\bm{M}$, and the classifier weights, $\bm{W}$. Our experiments will show that trojan attacks affect both the norm and angular properties of the features and classifier weights in different ways.
[…]
NC4: Nearest feature mean classification The inference time operation of the neural network reduces to a nearest neighbor decision rule in the final feature space. That is, a new test point is classified to the class with the feature mean nearest to its own feature representation. A metric for this is then the misclassification rate for this decision rule.
[…]
Extensive experiments conducted by [22] and [11] demonstrate that NC occurs across a wide variety of different architectures and datasets, with all of the the NC metrics consistently converging towards zero as models progress through the terminal phase of training. However, our observations indicate that application of a trojan trigger weakens this phenomenon. Specifically, under the same model settings and training processes, the NC1-4 metrics for trojaned models are frequently slower to converge than the benign models and often converge to a value larger those of analogous benign models.
(a) Equinorm metric [2] for the feature class means and the final classifier layer weight matrix.
(b) Equiangular metric [2] for the feature class means and the final classifier layer weight matrix.
3.3 Proposed Trojan Cleansing Method (ETF-FT)
Our experimental results in Section 4.1 strongly suggest that trojan attacks are incompatible with Neural Collapse. Unfortunately, it is difficult to use it to detect trojans–since the scale and rate of convergence of the Neural Collapse mechanisms varies strongly with the network architecture and dataset. Setting an expected degree of collapse associated with a trojan attack would be highly specialized to specific learning problems. Instead, we leverage our insights to design a method for cleansing a network of any potential trojan triggers, which we name ETF-FT.
ETF-FT aims to take a fully trained, possibly trojaned neural network and remove trojan triggers, if any exist, while maintaining performance on clean test data. As specified in Algorithm 1, we proceed by first over-writing the weights of the final fully-connected layer to a randomly generated simplex ETF, freezing those weights, and then fine-tuning the remaining model parameters on a small subset of clean data. Notably, this process works robustly with limited data, and the fine-tuning data does not necessarily need to come from the original training set.
Algorithm 1 Trojan Cleansing with Neural Collapse (ETF-FT)
Step 1: Replace weight matrix with $\textbf{W}_{\text{ETF}}$
Construct $\textbf{W}_{\text{ETF}}$ according to Algorithm 2
Replace the final layer weight matrix with $\textbf{W}_{\text{ETF}}$
Step 2: Fine-tune the model
Freeze the final layer weight matrix
Fine-tune on the clean dataset $\mathbf{X}_{\text{clean}}$
Output: Cleansed model $h_{\text{cleansed}}$
Algorithm 2 Constructing the $\mathbf{W}_{\text{ETF}}$ Matrix
Input: Number of classes $K$, random orthogonal matrix $\mathbf{P}\in\mathbb{R}^{m\times K}$
Initialize: $\mathbf{I}\leftarrow\mathbf{I}_{K}$triangleright Identity matrix of size $K\times K$
Initialize: $\mathbf{1}\leftarrow\mathbf{1}_{K\times K}$triangleright Matrix of ones of size $K\times K$
Compute $\mathbf{W}_{\text{ETF}}\leftarrow\sqrt{\frac{K}{K-1}}\times\mathbf{P}\times\left(\mathbf{I}-\frac{1}{K}\mathbf{1}\right)^{T}$
Output: Matrix $\mathbf{W}_{\text{ETF}}$
ETF-FT is motivated by our results which show that trojans work, in part, by deforming the symmetric structure NC describes in the weights of the final layer. By fixing those weights to be a random, but perfectly symmetric ETF we accomplish two things: first, we prevent the trojan deformation of the final layer and second, we reset the weights in a way known to be effective for classification. This addresses a challenge present in many trojan cleansing techniques: the more you perturb the poisoned model the less effective the trigger will be, but the harder is it to recover accuracy on clean data.
3.4 Experiment Details for Cleansing Methods Comparison
We compare our method with several state-of-the-art cleansing algorithms. Our evaluation considers two types of backdoor attacks: BadNets-A2O [10] and TrojanNN [17], and two model architectures: PreAct-ResNet18 [12] and ViT-B/16 [1].
We focus on two key metrics to evaluate the performance of each cleansing method: accuracy on clean data (ACC) and attack success sate (ASR). Accuracy reflects the model’s performance on clean data after cleansing, ideally showing no reduction compared to the original model. ASR, on the other hand, measures the percentage of poisoned samples classified into the attack target class. An effective cleansing method should reduce ASR to nearly zero.
We also evaluate the cleansing algorithm’s performance on a vision transformer, due to the popularity and efficacy of transformer architectures. However, many cleansing algorithms are not designed with transformers in mind. Where necessary, we made design choices, as in NPD, which inserts a new layer into the network, we chose to insert a linear layer before the final layer of the transformer encoders. We also did mild training hyperparameter tuning over batch size, learning rate, scheduler and decay rate for each algorithm including our own. Despite this, many algorithms show poor performance in this setting, relative to their results on ResNets. We want to emphasize that many of these algorithms are likely capable of better performance on transformers given careful adaptation. However, in the trojan attack scenario, the end user has outsourced network training to an adversary and is therefore unlikely to be a sophisticated deep learning engineer. As such, it is a significant strength of our algorithm that it can be applied to any classifier architecture without modification.
4 Results
4.1 Neural Collapse Disruption
@@
We first examine the behavior of the NC metrics described in Section 3.2 for both trojaned and benign models over the course of training on the CIFAR-10, CIFAR-100, and GTSRB datasets. Our results generally show that the application of a trojan trigger causes the NC metrics to converge more slowly and to converge to a larger value than their benign counterparts. There is variation in this behavior across different datasets and different metrics, but in every case the trojaned model shows weaker collapse in at least some metrics.
Figure 3 tracks NC2, defined in Equation 2, which characterizes the convergence of both the final classifier weights and the training feature means to a simplex ETF. The red dashed line in the figures marks the point where the training error reaches zero and so the regime in which Neural Collapse is generally observed. For the equinorm property, shown in Figure 3(a), the benign model shows the most uniformity in norm across both feature means and classifier weights, and as the percentage of trojaned data increases, the equinorm property weakens. This trend is particularly profound in the classifier weights. We note that, for both class means and classifiers, the source of this deviation is actually that the norm of the target class is consistently smaller than those of other classes. For the equiangular comparison, a similar, though smaller trend is observed. In Figure 3(b), the benign model reaches the smallest equiangular metric for both feature class means and classifiers. This further underscores the weakening of Neural Collapse due to trojaning.
NC3, defined in Equation 3, is shown in Figure 4. Here, again, the metric converges to a smaller value for the benign model than either trojaned model and the disruption to Neural Collapse appears to increase with the proportion of training data poisoned with the trojan trigger.
Figure 4: Self-duality Comparison of the convergence of trojaned and benign models on the NC3 metric [3] for CIFAR-10 on ResNet18.
But this pattern is not observed in every collapse metric for every dataset. On CIFAR-10 we did not observe a significant difference between the benign and trojaned models for NC1, characterizing the variability across feature representations, or NC4, characterizing how close the network’s decision rule is to a nearest neighbor rule in the final feature space, as they all go down to zero. However, collapse is noticeably weakened in these metrics in our experiments on CIFAR-100 and GTSRB, as shown in Figures 6 and 6 and Table 1.
@@
Figure 5: Variability Collapse Comparison of the convergence of trojaned and benign models on the NC1 metric [1] for GTSRB on ResNet18.
@@
Figure 6: Nearest feature classification Comparison of the convergence of trojaned and benign models on the NC4 metric [4] for GTSRB on ResNet18.

### v2

@@ 1 Introduction
Over the past decade, deep neural networks have achieved state-of-the-art performance in a vast array of sensitive machine learning (ML) tasks such as autonomous driving, medical diagnostics, and financial portfolio management. The unprecedented speed of this adoption, along with prior research establishing the vulnerability of these methods to a wide variety of attacks, has made the study of security and safety of ML models an important area of research. This paper considers an important class of training-time attacks on these models known as trojan attacks.
Trojan attacks, or backdoor attacks, aim to maliciously corrupt a neural network’s training process such that the presence of a trojan trigger in any input will force the network to produce a target classification [10]. This type of attack is generally accomplished by means of data poisoning, where the adversary adds their chosen trigger to a small portion of a training dataset and labels the poisoned samples with their selected target class. Models trained on this dataset then learn both the correct classifications for clean data and the malicious classification for inputs containing the trigger. Fig. 1, provided in [27], illustrates an example of clean data point (a) along a poisoned one (b). The yellow trigger in this example is clearly visible, but [25] showed that poisoning may take on a subtle nature that is undetectable by human auditors so that that malicious behavior can be hidden until test time or model deployment.
Figure 1: An example of trojan attack on a traffic sign classifier. (a) model works normally on clean data without triggers, (b) attackers manipulate the prediction by adding a specific trigger. Image due to [27].
In this paper, for the first time, we make a concrete connection between training with poisoned data and a phenomenon recently studied in the deep learning theory literature, called Neural Collapse [24]. Neural Collapse describes the widely observed phenomenon wherein, over the course of training a neural network, highly symmetric structures emerge in both the network’s weights and the feature representations of training samples. In contrast, trojan attacks are inherently asymmetric: they require any input be classified to a single target class given only the presence of a small perturbation. We hypothesize and experimentally demonstrate that the injection of a trojan trigger into a neural network breaks the symmetries of Neural Collapse. We then use these results to develop a carefully targeted, lightweight method for mitigating trojan attacks that is both architecture and dataset agnostic.
@@ 1.1 Contributions
There is a large body of work on trojan detection and mitigation, detailed in Sec. 2.2. Some of these approaches first attempt to understand how trojan triggers alter the structure of neural networks: Neural Cleanse [29] suggests that triggers activate a certain subset of neurons, the topological analysis of [35] finds that triggers introduce an unusual number of skip connections between early and late layers in the network. But these analyses are generally unable to predict precisely where and how a trigger will manifest in a particular neural network and must rely on data-driven approaches to identify, for instance, the particular subset of neurons which comprise the action of the trigger.
We pursue this question through the inherent asymmetry introduced by training on poisoned datasets and the natural impact this has on the Neural Collapse (NC) phenomenon. We discuss NC in depth in Sec. 2.3, but, briefly, NC is a feature of training over-parameterized models where, as shown in Fig. 2, the final feature representations of the training data and the weight matrix of the final layer of the network, both, converge to a simple and highly symmetric geometric structure. We hypothesize that this symmetric structure, however, directly contradicts the asymmetry introduced by data poisoning. In this work, we present extensive experimental evidence supporting this hypothesis and show that the process of embedding a trojan in a neural network significantly weakens the NC phenomenon in measurable and predictive ways.
We leverage these insights to initiate a new class of trojan cleansing methods: algorithms which take a potentially trojaned network and by imposing NC-inspired symmetry remove the trojan trigger while maintaining the network’s performance on clean data. In other words, identifying the predictive impact the trojan embedding has on NC metrics allows us to arrive at an entirely new method for cleansing. Many existing cleansing methods, as detailed in Sec. 2.2, rely either on some form of model compression, which may harm the model’s generalization capabilities, or on reconstructing the trojan trigger, which relies on knowledge of the class of the triggers the adversary may have chosen from. In contrast, we are able to introduce a small and carefully targeted adjustment to a small subset of the weights of the network without any knowledge of the possible trojan triggers: mitigating the attack without deteriorating the generalization power of the model.
@@ 2 Related Work
Figure 2: The figure illustrates NC as training progresses from left to right. Green spheres represent the vertices of an ETF, red points represent the linear classifiers, large blue/green points represent class means, and small blue/green points represent final layer features, with different shades indicating different classes. As training progresses, the features concentrate. Figure due to [24].
@@ 2.2 Model Cleansing
Li et al. [17] and Xia et al. [31] use forms of distillation to preserve only the model’s performance on clean data while Wang et al. [29] identifies and prunes neurons likely associated with the trojan trigger. Other approaches rely on first reconstructing potential triggers and then attempt to remove them: [4] learns a generative model to create likely triggers and then corrects the network via adversarial training. Zeng et al. [33] proposes a minimax formulation and computationally efficient optimization scheme for trigger reconstruction. Other approaches include [15], which leverages sparse reconstruction of inputs and feature representations to remove potential triggers, and [37], which applies adversarial training to a new layer inserted into the network. [36] is a detection and mitigation method that shrinks the norms of backdoor-related neurons by using sharpness-aware minimization with fine-tuning. Wu et al. [30] introduced BackdoorBench, a platform which implements many of these algorithms and provides a leader board that guides our benchmark selection in Sec. 4.2.
We note that our work is related but fundamentally distinct from trojan detection which aims to determine whether a model may be trojaned, including those focusd on data analysis, as in Baracaldo et al. [2], Liu et al. [18], Shen et al. [26] as well as those work that directly analyze the model, such as Chen et al. [3], Ma et al. [21], Fields et al. [9].
@@ 2.3 Neural Collapse
Neural Collapse is a phenomenon, first reported by Papyan et al. [24] and thoroughly reviewed by [16], which has been observed across a wide variety datasets and model architectures. It describes a process observed in over-parameterized classification deep neural networks which are trained past zero training error, into a period referred to as the terminal phase of training (TPT). As depicted in Fig. 2, as the TPT progresses, the final feature representations of the training data and the weight matrix of the network’s last layer converge to a highly uniform, symmetric geometric structure known as a simplex equiangular tight frame (simplex ETF).
While NC is associated with models trained past perfect training accuracy, this practice is known to have practical benefits as in the double descent phenomenon [22]. Further experiments in [24] demonstrate in particular that the TPT offers advantages including improved generalization performance, enhanced robustness, and increased interpretability. Various theoretical models, based on locally elastic dynamics [34] and unconstrained features [38, 37] have begun examining how NC arises and its connections to these other neural network characteristics.
It is important to note that Neural Collapse is not a binary phenomenon that either does or does not occur. In practice the characteristics of collapse arise even before the TPT and gradually converge over the course of the TPT. Consequently, our discussion of NC will focus on its degree, quantified by four metrics which converge to zero as the qualitative characteristics of NC obtain. We will define these metrics in Sec. 3.2 after introducing notation.
3 Neural Collapse Disruption Hypothesis
[24] and [12] observed and proposed metrics for Neural Collapse across diverse architectures and datasets. In this section, we review these metrics, and provide evidence that trojan triggers weaken Neural Collapse phenomenon. In particular, we evaluate various NC metrics associated with a variety of trojaned models and show that every trojaned model has at least one weaker NC metric. Furthermore, benign models consistently achieve at least the same degree of collapse as their trojaned counterparts across all metrics.
3.1 Neural Collapse: Definitions
@@
The trojan attacks we evaluate in our experiments are carried out via data poisoning. To implement such an attack, we first choose a trigger function $\kappa:\mathbb{R}^{d}\to\mathbb{R}^{d}$. In our experiments this function will be the overlay of a small, solid patch on the base image, as shown in Fig. 1, or the application of an Instagram filter to the base image, as shown in Fig. 7. We then choose a poisoning proportion $\delta\in[0,1)$ and a target class $k_{T}\in[K]$. For each $\bm{X}_{k}$ with $k\neq k_{T}$ and each sample $\bm{x}\in\bm{X}_{k}$, with probability $\delta$ we delete $\bm{x}$ from $\bm{X}_{k}$ and add $\kappa(\bm{x})$ to $\bm{X}_{k_{T}}$. We say a model is trojaned if $\delta>0$, and a model is benign if $\delta=0$.
@@ 3.2 Neural Collapse: Metrics and Trojan Impact
The observed phenomenon of Neural Collapse can be more precisely described by a set of four characteristics laid out in [24] and quantified by a standard set of metrics defined in [12] and used in other analyses of NC such as [11] and [14]. In this section we will describe these metrics and present results comparing their convergence over the course of training for both benign and trojaned models on CIFAR-10, CIFAR-100, and GTSRB datasets. Our results show that the application of a trojan trigger causes the NC metrics to converge more slowly and to a larger value than their benign counterparts. There is variation in this behavior across different datasets and different metrics, but in every case the trojaned model shows weaker collapse.
@@
Figure 3: Variability Collapse Comparison of the convergence of trojaned and benign models on the NC1 metric [Eq. 1] for GTSRB on ResNet18.
@@
Figure 4: Self-duality Comparison of the convergence of trojaned and benign models on the NC3 metric [Eq. 3] for CIFAR-10 on ResNet18.
Figure 5: Nearest feature classification Comparison of the convergence of trojaned and benign models on the NC4 metric [Eq. 4] for GTSRB on ResNet18.
NC1: Variability collapse The final feature representation of every training sample in every class converges to the mean of the feature representations of the samples from their respective classes. Equivalently, the variability of feature representations within any given class vanishes. This can be quantified by defining the within-class and between-class covariances
$\Sigma_{B}$ $\coloneqq\frac{1}{K}\sum_{k=1}^{K}\bm{m}_{k}\bm{m}_{k}^{T}.$
The degree of variability collapse is then quantified by the relationship between the spread of features within a class relative to the spread between the feature means of different classes:
[…]
Experimental evaluation of NC1 for both benign and trojaned models are shown in Fig. 5 and Tab. 1. On CIFAR-10 we did not observe a significant difference between the benign and trojaned models for this metric, but convergence of NC1 was significantly weaker in our experiments on CIFAR-100 and GTSRB. Note that the red dashed line in all figures represents the point at which the models achieve zero training error, and so the beginning of the terminal phase of training.
NC2: Convergence to a simplex ETF Both $\bm{M}$ and $\bm{W}$ converge to a simplex ETF. In particular, this means that the norms of each of the rows equalize and that they become maximally pairwise separated from each other as measured by cosine similarity. Note that the maximum pairwise cosine similarity among a set of $K$ vectors is $(1-K)^{-1}$. These two characteristics can be quantified, respectively, by the two statistics
[…]
(a) Equinorm metric [Eq. 2] for the feature class means and the final classifier layer weight matrix.
(b) Equiangular metric [Eq. 2] for the feature class means and the final classifier layer weight matrix.
Fig. 6 tracks Eq. 2, which characterizes the convergence of both the classifier weights, $\bm{W}$ and the centered class means, $\bm{M}$ to a simplex ETF. For the equinorm property, shown in Fig. 6(a), the benign model shows the most uniformity in norm across both feature means and classifier weights, and as the percentage of trojaned data increases, the equinorm property weakens. This trend is particularly profound in the classifier weights. We note that, for both class means and classifiers, this is caused by the fact that the norm of the target class is consistently smaller than those of other classes. For the equiangular property a similar, though smaller trend is observed.
[…]
Fig. 5 shows this metric also converges to a smaller value for the benign model than either trojaned model and the disruption to Neural Collapse increases with the proportion of training data poisoned.
NC4: Nearest feature mean classification At inference time, the neural network uses a nearest neighbor decision rule in the final feature space. That is, a new test point is classified to the class with the feature mean nearest to its own feature representation. A metric for this is then the misclassification rate for this decision rule.
[…]
Similar to NC1, on CIFAR-10 we did not observe a significant difference between the benign and trojaned models for NC4 as the nearest neighbor rule reaches perfect accuracy for each of them. But collapse is noticeably weakened in this metric in our experiments on CIFAR-100 and GTSRB, as shown in Fig. 5 and Tab. 1.
4 Trojan Cleansing Method (ETF-FT)
4.1 Methodology
This tension motivates an algorithm for cleansing a network of any potential trojan triggers, which we name ETF-FT. ETF-FT aims to take a fully trained, possibly trojaned neural network and remove trojan triggers, if any exist, while maintaining the network’s good performance on clean test data. The effects of a trojan attack are most apparent in Fig. 6(a), which shows that the weights of the final layer of a benign neural network exhibit a much greater degree of symmetry than those of an analogous trojaned network. So, as specified in Algorithm 1, we proceed by first over-writing the weights of the final fully-connected layer to a randomly generated simplex ETF, freezing those weights, and then fine-tuning the remaining model parameters on a small subset of clean data. Notably, this process works robustly with limited data, and the fine-tuning data does not necessarily need to come from the original training set.